Y+4x=4 and 6c+2y=6
Answer by Simnepi(216)

(I assume the 'c' in the second equation is supposed to be an 'x'!)
y+4x=4 Equation (1)
6x+2y=6 Equation (2)
From equation (1) we can see that
y=4-4x Equation (3)
Now we substitute this value we have for y into equation (2)
So now we have
6x+2*(4-4x)=6
Multiplying the brackets gives
6x+8-8x=6
simplifying this gives
-2x=-2
therefore
x=1
now we can use this to find out y
substitute this value for x into either (1), (2) or even (3)!
Let's use (1)
y+4*1 = 4
we now find that
y = 0